# Defect Positioning in Combinatorial Metamaterials

### v1

@@ I Introduction
Combinatorial metamaterials allow to design which loops will be compatible and which frustrated. Thus they are useful for exploring the effects of frustration and defects in physical systems in general Alexander et al. (2012). Mechanical compatibility leads to combinatorial matching rules between neighboring units, with similar concepts employed in origami Dieleman et al. (2020). Manipulating the discrete orientations of the building blocks allows to design metamaterials with textured response Coulais et al. (2016) or to use topological defects for spatially steering stress fields Meeussen et al. (2020); Pisanty et al. (2021). Computational methods have been employed to obtain multiple desired responses of a given metamaterial when there are two soft modes per building block Bossart et al. (2021); van Mastrigt et al. (2022, 2023).
@@
In Sec. IV we investigate the global structure of mechanical incompatibilities, or defects, where there is frustration between the desired deformations of neighboring blocks. We study the possible spatial distributions of point defects for the different 2D blocks, and the possible topological knotted or linked structures that defect lines can form for the various 3D blocks. Both in 2D and in 3D, we identify block types for which an arbitrary distribution of defects may not necessarily be realizable. However, in 3D we show that defect lines with arbitrary knottedness may be constructed using any one of the non-trivial blocks. Finally, in Sec. V we discuss the results, and their consequences for further work on characterizing and designing physical responses of combinatorial metamaterials.
@@ III.1 Local condition for compatibility
We will test for compatibility locally. Namely, for each lattice, we will consider all the minimal loops within it. In 2D, a minimal loop is a loop containing the blocks surrounding a vertex in the lattice, and in 3D it contains the blocks surrounding an edge in the lattice. Any larger loop in the lattice may be constructed from a combination of such minimal loops, and therefore, if all the minimal loops are compatible, so is any larger loop that they form Meeussen et al. (2020).
In Fig. 2 we examine mechanical compatibility for a minimal loop in the square lattice, which includes four building blocks around a vertex in the lattice. Here, we use Block S3 to demonstrate that certain orientations of the blocks give a compatible structure, while other orientations lead to frustration. If a minimal loop in 2D is frustrated, we can identify the vertex that this loop encircles as a mechanical defect in the lattice. In 3D, edges will be the basic objects that are either compatible or frustrated (a.k.a. defected).
Now in all cases, whether a vertex in 2D or edge in 3D is defected depends solely on the parity of the number of hinges that are incident to it: Put simply, odd means defected, and even is non-defected. Indeed, as we traverse our minimal loop, the rotational direction of the incident facet flips with every hinge, and the condition for consistency is exactly that the number of such flips be even.
Figure 2: Mechanical compatibility. Four Block S3 square units can be mutually oriented, (a) in a compatible manner such that they can all simultaneously deform in their soft mode, or (b) in a frustrated manner in which there must be at least one block that cannot deform in its soft mode. The number of hinges incident to the central vertex is two (even) in (a) and three (odd) in (b). The red dot indicates the defected vertex at the center of the frustrated loop.
@@ III.2 Defects form closed curves
First, consider a single cubic block. Its 2D surface is divided into regions of two different colors, indicating inward and outward motion in the soft mode, as depicted in Fig. 1b. The one-dimensional boundary of each color is a cycle, i.e., it has even degree at each vertex. Since two adjacent faces connected by a hinge have the same color and two faces connected with a strut have opposite colors, it follows that there is an even number of struts at each vertex. Now, since each vertex of the cubic block has total degree three, the number of hinges at each vertex is odd.
When considering the cubic lattice, each edge is shared by four blocks. Each of these four blocks has either a strut or a hinge at that edge, and the edge is defected if the number of hinges is odd. The parity of defected edges meeting at a vertex of the lattice is the parity of the total number of these hinges. This number may be considered by summing over the six edges, where for each edge we sum over the four blocks surrounding it. Alternatively, we can sum over the eight blocks around the vertex, noting that each such block has three edges, where as discussed above, the number of hinges among the three must be odd. Since the sum of eight odd numbers is even, the number of defected edges meeting at the vertex must be even.
@@ IV.3.2 Block C6 defect realization
Since in Block C6 all vertices have exactly one edge with a hinge, this simple scanning method cannot be directly applied to it, and a more delicate procedure is required. Here we provide a construction protocol for realizing with Block C6 any defect configuration, which satisfies the parity constraint explained in Sec. III above. Namely, the number of defected edges meeting at any vertex inside the lattice must be even. A useful practical consequence of this is that if we design the metamaterial while making sure that the defectedness of five of the six local edges is as in the desired defect configuration, then the parity constraint will assure that the sixth edge fits the prescription, too.
the upper side of any white square is not needy. (*)
The trick will be to maintain this property for all subsequent rows, too. (Here, ‘top’ and ‘upper’ are meant in the sense of the $y$ coordinate.) Note that as we proceed from top to bottom, the set of needy edges will be continuously updated, and (*IV.3.2) is meant with respect to the set of needy edges just before we start working on a new row.
Indeed as long as (*IV.3.2) holds, it makes sense to carry out the following steps. For each yellow square along our row, we check which of its horizontal sides — the edges in the $x$ direction — are needy. We assign the orientation $zy$ if the answer is none; $zx$ if the answer is both; $yx$ if it is only the top; and $yz$ if it is only the bottom. Then, we look at the white squares along the same row and consider their vertical sides — the edges in the $y$ direction. If none is needy, we assign orientation $zy$ for the block which is responsible for this white square; if both are needy, then $yz$; if only the left or the right side is needy, then $xy$ or $xz$, respectively. This takes care of vertical edges along this row. Furthermore, because of (*IV.3.2), needy edges do not remain along the upper boundary, either.
Finally, as to the lower boundary, the construction ensures that only the bottom edges of white squares may remain (or become) needy. In terms of the next row, these are top edges of yellow squares. In other words, the upper sides of the white squares of the next row are not needy, that is to say (*IV.3.2) holds for the next row of squares and the iteration may continue unabated in the negative $y$ direction. After that, there is no impediment to continuing the process in the negative $z$ direction, either, until the construction of the metamaterial is complete.
@@ IV.3.3 Block C2 counting argument for non-realizability
In order to find a defect set which is non-realizable with Block C2, we wrote a computer program that maps our problem onto an instance of the Boolean satisfiability problem, also known as SAT, to check if a given set is realizable or not def ; see Appendix B for more details. To generate candidate defect sets, whose realizability the program would check, we first solved a forward problem of finding the defects in a randomly chosen structure of Blocks C3.
@@ IV.3.4 Block C2 arbitrary knottedness
In order to build the over- and under-crossings, we make some further orientation changes. Namely, we re-assign the orientations of the same set of blocks we started with (corresponding to the shaded squares), but this time at the middle ($z=0$) level in such a way that each time, the boundaries of two squares, perpendicular to the $x$-axis this time, are added to the defect set, cf. Fig. 8b. In Fig. 8c we show an example of the outcome. In Appendix C we give specific block assignments which illustrate our method.
@@ V Discussion
The effect of defects on the mechanical response of metamaterials has been demonstrated Meeussen et al. (2020); Pisanty et al. (2021). It would be interesting to understand how the presence of multiple defects, possibly with a complex spatial distribution, affects the mechanical response. Another direction in this regard would be identifying mechanical consequences stemming from the topology (as in, knottedness) of defect lines within metamaterials, which we are now able to to generate. It would also be interesting to extend our approach of generating defect lines with arbitrary knottedness to situations with defect line crossings, where vertices have 4 or 6 defected edges.
Finally, for the 3D blocks that we considered, defect lines cannot terminate or split in the bulk. It would be interesting to identify metamaterials that could have point defects in 3D, similarly to hedgehog point defects in nematic liquid crystals.
@@ Appendix A H6 arbitrary defect positioning
Appendix B Defects realization program
@@
Appendix C Concrete example of a knotted defect

### v2

@@ I Introduction
Combinatorial metamaterials allow to design (with certain limitations that we discuss below) which loops will be compatible and which frustrated. Thus they are useful for exploring the effects of frustration and defects in physical systems in general Alexander et al. (2012). Mechanical compatibility leads to combinatorial matching rules between neighboring units, with similar concepts employed in origami Dieleman et al. (2020). Manipulating the discrete orientations of the building blocks allows to design metamaterials with textured response Coulais et al. (2016) or to use topological defects for spatially steering stress fields Meeussen et al. (2020); Pisanty et al. (2021). Computational methods have been employed to obtain multiple desired responses of a given metamaterial when there are two soft modes per building block Bossart et al. (2021); van Mastrigt et al. (2022, 2023).
@@
In Sec. IV we investigate the global structure of mechanical incompatibilities, or defects, where there is frustration between the desired deformations of neighboring blocks. We study the possible spatial distributions of point defects for the different 2D blocks, and the possible topological knotted or linked structures that defect lines can form for the various 3D blocks. Both in 2D and in 3D, we identify block types for which an arbitrary distribution of defects is not realizable. However, in 3D we show that defect lines with arbitrary knottedness may be constructed using any one of the non-trivial blocks. Finally, in Sec. V we discuss the results, and their consequences for further work on characterizing and designing physical responses of combinatorial metamaterials.
@@ III.1 Local condition for compatibility
We will test for compatibility locally. Namely, for each lattice, we will consider all the minimal loops within it. In 2D, a minimal loop is a loop containing the blocks surrounding a vertex in the lattice, and in 3D it contains the blocks surrounding an edge in the lattice. Any larger loop in the lattice may be constructed from a combination of such minimal loops, and therefore, if all the minimal loops are compatible, so is any larger loop that they form Meeussen et al. (2020). More precisely, let us emphasize that in this paper we only consider metamaterials enclosed in simply connected regions; in those cases, in the sense of mod $2$ chains, any loop can indeed be written as a sum of elementary loops. Furthermore, if we associate 00 to compatible loops and $1$ to incompatible ones, we obtain a mod $2$ valued mapping that is additive on all loops.
Figure 2: Mechanical compatibility. Four Block S3 square units can be mutually oriented, (a) in a compatible manner such that they can all simultaneously deform in their soft mode, or (b) in a frustrated manner in which there must be at least one block that cannot deform in its soft mode. The number of hinges incident to the central vertex is two (even) in (a) and three (odd) in (b). The red dot indicates the defected vertex at the center of the frustrated loop. (c) Eight Block C4 cubic units oriented in a compatible manner; all six edges have an even number of hinges (blue lines). (d) A frustrated orientation of these eight units with two defected edges (red lines), at which there is an odd number of hinges.
In Fig. 2a,b we examine mechanical compatibility for a minimal loop in the square lattice, which includes four building blocks around a vertex in the lattice. Here, we use Block S3 to demonstrate that certain orientations of the blocks give a compatible structure, while other orientations lead to frustration. If a minimal loop in 2D is frustrated, we can identify the vertex that this loop encircles as a mechanical defect in the lattice. In 3D, edges will be the basic objects that are either compatible or frustrated (a.k.a. defected).
In 2D, we use the following formalism to determine if a vertex is defected or not. We start with an arbitrary facet surrounding this vertex and associate to it $S_{0}=1$ to indicate that it wishes to deform clockwise around the vertex. As we move to the next adjacent facet, if the corner of the block separating the two facets includes a strut, then that next facet will prefer to deform in the same orientation, that is $S_{1}=S_{0}=1$, while if the block contains a hinge, the direction of rotation around the vertex will flip, resulting in $S_{1}=-S_{0}=-1$. As we encircle the vertex and return to the initial facet, a compatible deformation of all blocks will, therefore, require this facet to deform in the rotational direction described by $S=(-1)^{h}$, where $h$ is the number of hinges incident to the vertex. Clearly, if the number of hinges is even, this is equal to $S_{0}=1$ and the blocks around this vertex can simultaneously deform to their soft mode, while if $h$ is odd, the vertex is frustrated and we identify it as a mechanical defect. A similar analysis can be carried out in 3D, where we consider the blocks surrounding an edge of the lattice, and conclude that the defectedness of that edge depends on the parity of hinges around it.
@@ III.2 Defects form closed curves
First, consider a single cubic block. Its 2D surface is divided into regions of two different colors, indicating inward and outward motion in the soft mode, as depicted in Fig. 1b. Adjacent faces of the same color are separated by a hinge, and a pair of opposite colors are separated by a strut. The one-dimensional boundary of each color is a cycle, i.e., it has even degree at each vertex. Since two adjacent faces connected by a hinge have the same color and two faces connected with a strut have opposite colors, it follows that there is an even number of struts at each vertex. Now, since each vertex of the cubic block has total degree three, the number of hinges at each vertex is odd.
When considering the cubic lattice, each edge is shared by four blocks, as shown in Fig. 2c,d. Each of these four blocks has either a strut or a hinge at that edge, and the edge is defected if the number of hinges is odd. The parity of defected edges meeting at a vertex of the lattice is the parity of the total number of these hinges. This number may be considered by summing over the six edges, where for each edge we sum over the four blocks surrounding it. Alternatively, we can sum over the eight blocks around the vertex, noting that each such block has three edges, where as discussed above, the number of hinges among the three must be odd. Since the sum of eight odd numbers is even, the number of defected edges meeting at the vertex must be even.
@@ IV.3.2 Block C6 defect realization
Since in Block C6 all vertices have exactly one edge with a hinge, this simple scanning method cannot be directly applied to it, and a more delicate procedure is required. In Appendix B we provide a construction protocol for realizing with Block C6 any defect configuration which satisfies the parity constraint explained in Sec. III above.
@@ IV.3.3 Block C2 counting argument for non-realizability
In order to find a defect set which is non-realizable with Block C2, we wrote a computer program that maps our problem onto an instance of the Boolean satisfiability problem, also known as SAT, to check if a given set is realizable or not def ; see Appendix C for more details. To generate candidate defect sets, whose realizability the program would check, we first solved a forward problem of finding the defects in a randomly chosen structure of Blocks C3.
@@ IV.3.4 Block C2 arbitrary knottedness
In order to build the over- and under-crossings, we make some further orientation changes. Namely, we re-assign the orientations of the same set of blocks we started with (corresponding to the shaded squares), but this time at the middle ($z=0$) level in such a way that each time, the boundaries of two squares, perpendicular to the $x$-axis this time, are added to the defect set, cf. Fig. 7b. In Fig. 7c we show an example of the outcome. In Appendix D we give specific block assignments which illustrate our method.
@@ V Discussion
The effect of defects on the mechanical response of metamaterials has been demonstrated, for instance by stressed regions bypassing topological defects in 2D Meeussen et al. (2020). It would be interesting to understand how the presence of multiple defects, possibly with a complex spatial distribution, affects the mechanical response. Realization of metamaterials with non-trivial defect structures, and specifically with complex knotted topologies, is interesting in its own right. Another direction in this regard would be identifying mechanical consequences stemming from the topology (as in, knottedness) of defect lines within metamaterials, which we are now able to to generate. Closed loops of defects can cause stress to concentrate in the region enclosed by the loop, or alternatively to avoid this region Pisanty et al. (2021). We are not currently aware of specific mechanical implications for realizing knotted defect lines, however we note that knotted floppy chains have been recently realized in combinatorial metamaterials Liu et al. (2025), and suggest that it would be interesting to explore, for instance, how linking several defect loops alters stress steering in metamaterials.
It would also be interesting to extend our approach of generating defect lines with arbitrary knottedness to situations with defect line crossings, where vertices have 4 or 6 defected edges. Another intriguing possibility would be dynamically repositioning defects — for example, by using bistable blocks — in ways that would preserve topological features. Finally, for the 3D blocks that we considered, defect lines cannot terminate or split in the bulk. It would be interesting to identify metamaterials that could have point defects in 3D, similarly to hedgehog point defects in nematic liquid crystals.
@@ Appendix A H6 arbitrary defect positioning
Appendix B Defects realization with Block C6
Since in Block C6 all vertices have exactly one edge with a hinge, a simple scanning method cannot be directly applied to it, and a more delicate procedure is required. Here we provide a construction protocol for realizing with Block C6 any defect configuration, provided that it satisfies the parity constraint explained in Sec. III in the main text. Namely, the number of defected edges meeting at any vertex inside the lattice must be even. A useful practical consequence of this is that if we design the metamaterial while making sure that the defectedness of five of the six local edges is as in the desired defect configuration, then the parity constraint will assure that the sixth edge fits the prescription, too.
the upper side of any white square is not needy. (*)
The trick will be to maintain this property for all subsequent rows, too. (Here, ‘top’ and ‘upper’ are meant in the sense of the $y$ coordinate.) Note that as we proceed from top to bottom, the set of needy edges will be continuously updated, and (*B) is meant with respect to the set of needy edges just before we start working on a new row.
Indeed as long as (*B) holds, it makes sense to carry out the following steps. For each yellow square along our row, we check which of its horizontal sides — the edges in the $x$ direction — are needy. We assign the orientation $zy$ if the answer is none; $zx$ if the answer is both; $yx$ if it is only the top; and $yz$ if it is only the bottom. Then, we look at the white squares along the same row and consider their vertical sides — the edges in the $y$ direction. If none is needy, we assign orientation $zy$ for the block which is responsible for this white square; if both are needy, then $yz$; if only the left or the right side is needy, then $xy$ or $xz$, respectively. This takes care of vertical edges along this row. Furthermore, because of (*B), needy edges do not remain along the upper boundary, either.
Finally, as to the lower boundary, the construction ensures that only the bottom edges of white squares may remain (or become) needy. In terms of the next row, these are top edges of yellow squares. In other words, the upper sides of the white squares of the next row are not needy, that is to say (*B) holds for the next row of squares and the iteration may continue unabated in the negative $y$ direction. After that, there is no impediment to continuing the process in the negative $z$ direction, either, until the construction of the metamaterial is complete.
Appendix C Defects realization program
@@
Appendix D Concrete example of a knotted defect